OpenStudy (anonymous):

10.23 as a mixed number

OpenStudy (debbieg):

Writing as a mixed number simply means to right as the whole number part and the fraction part. For example, if I wanted to write 5.6 as a mixed number, I see that the whole number part is 5, and he fraction part is what is after the decimal: 6 tenths. So I write that as a fraction: \[\Large 5+\frac{ 6 }{ 10 }\]and then I reduce and write them "together", as is the format for a mixed number: \[\Large 5+\frac{ 6 }{ 10 }= 5+\frac{ 3 }{ 5} = 5\frac{ 3 }{ 5}\] Now can you apply that example to 10.23?